OpenStudy (anonymous):

how do I solve... Identify the relative minimum value for the function shown below. G(x)=x^3-3x^2+2

OpenStudy (zale101):

Calculus problem? If so, take the derivative of G(x)=x^3-3x^2+2 and then set it equal to zero to find the critical points.

OpenStudy (solomonzelman):

you can either take the derivative, set it equal to zero, and find all the critical numbers. Then plug in these critical numbers into the function, and the smallest output would be the absolute minimum. Or you can graph it, and visualize it like that.

OpenStudy (solomonzelman):

@lexismith is this calculus, or not?

OpenStudy (anonymous):

Pre-Calc. yes

OpenStudy (solomonzelman):

Also, one more thing. You haven't given the interval to us

OpenStudy (solomonzelman):

on what interval you are taking about? if the interval is \(\large\color{blue}{ (-\infty,+\infty) }\) then the minimum is \(\large\color{blue}{ -\infty }\). What are the boundaries?

OpenStudy (solomonzelman):

the question asks to find the minimum value of \(\large\color{blue}{ G(x)=x^3-3x^2+2 }\), over the interval of?? (it must be giving you the interval)

OpenStudy (solomonzelman):

he's offline-:(

OpenStudy (zale101):

:(

OpenStudy (p0sitr0n):

that feel when you took time to type the answer and the asker doesnt need it anymore tho